Attendees: Bruce Glagola, Dean Wyncott, Nena Moonier, Wendy VanWingeren, Clay White, Bruno Fieramosca, and Patti Pedergnana
- Tech Tuesday has resumed. Dean will attend the 438 meeting, Clay the 433 meeting, and Bruno the 431 meeting.
- Bruce asked Clay to update all the shop access lists. Clay is walking through all of the machine shops with Elroy today.
- Joe Gagliano was notified that not all BIVs were open for start up.
- In the next shutdown, 6-ID will get a new Front End and 4-ID's vacuum chamber will be replaced.
- 9-BM's monochromator is mostly in. The beamline should start up some time next week. The beamline personnel are aware that a HP type C survey is needed.
- There will be a BSL-2 experiment running at 2-ID next week. There is a brain sample coming on site the week after that.
- MOM is starting their door maintenance at 1-BM today and may add on 1-ID's door maintenance if there is time.
- Sam Weir is running radioactive samples at 16-BM-D starting Wednesday, February 17th.
- Scheck is working on their first door repair at 16-BM-D today.
- 11-ID-C had an ODH monitor fail over the weekend. Bruno is going to replace the sensor on the unit.
- There was a 911 on Sector 2 beamline personnel this morning.
- CVD Equipment will do a deposition laboratory installation soon.
- 1-ID has a new Remote Shutter Interface.
- Paul Rossi recommends that we change our ODH monitors to the NSLSII style monitors and phase them in over five years time. The cost will be about $65K.
- There is a new application to order kitchen appliances. Bruno has an order in for a new microwave now.
- 4-ID's mini-hutch is open allowing Steve Davey and Julie Cross to view the K5 collimator. There will be a water flow adjustment on the collimator. Bruno will post two CCWPs for this work: one for opening the mini-hutch and one for the water adjustment.